OpenStudy (anonymous):

f(x) = 3x – 7 and g(x) = –2x – 6. Find (ƒ ∘ g)(4).

OpenStudy (anonymous):

\[f\circ g(4)=f(g(4))\] so first find \(g(4)=-2\times 4-6=-8-6=-14\)

OpenStudy (anonymous):

then \(f(g(4))=f(-14)=3\times (-14)-7\) whatever that is

OpenStudy (hba):

Do you understand ? @HorrificallyDelicious

OpenStudy (anonymous):

the first part i have -14 how do you get g(x)

OpenStudy (anonymous):

i have -70 as my answer, by -14(5)

OpenStudy (anonymous):

you have \(g(4)=-14\) so now your job is to find \(f(-14)\)

OpenStudy (anonymous):

\(f(-14)=3\times (-14)-7\) and you need to compute that number

OpenStudy (anonymous):

right ans -49

OpenStudy (anonymous):

how did you get -49, i got -70

OpenStudy (hba):

Because your'e doing it wrong.

OpenStudy (hba):

We have to find f(-14) so 3(-14)+7=?

OpenStudy (hba):

Do it again

OpenStudy (hba):

3(-14)-7=-42-7=-49 got it ?

OpenStudy (anonymous):

yes thanks

OpenStudy (hba):

Your'e welcome intelligent person.